# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ 3. Continual Mechanisms
By Definition 2.12, an interactive mechanism $\mathcal{M}$ with initial state space $S_{\mathcal{M}}^{\text{init}}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$, if for any two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and any adversary $\mathcal{A}$, the view of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ is $(\epsilon,\delta)$-indistinguishable. This means that for the worst-case (i.e., adversarial) choice of neighboring initial states, an adversary by adaptively asking queries should not be able to determine whether $\mathcal{M}$ is running on $s_{0}$ or $s_{1}$.
@@
To formalize this, we restrict attention to adversaries that send a pair of neighboring initial states $(s_{0},s_{1})$ as their first message and pairs of identical queries $(q_{i},q_{i})$ in its subsequent messages. We also define an IPM $\mathcal{I}$ with the initial state space $S_{\mathcal{I}}^{\text{init}}=\{0,1\}$ to stand between the adversary and $\mathcal{M}$, filtering the pairs based on its, i.e., $\mathcal{I}$’s, initial state. Specifically, for $b\in\{0,1\}$, in an interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$. With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for any adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
@@
To capture ”event-level privacy” (where the privacy unit is a single record), Dwork et al. [7] define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in a single step. They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for any two neighboring input sequences. This definition does not allow for data sequences that are adaptively chosen based on the answers released.
@@
A mechanism $\mathcal{M}$ in the continual observation model can be viewed as an IM with a single possible initial state, representing an empty dataset. The input message set of $\mathcal{M}$ consists of the empty record and all possible data records. Instead of having a challenge step, we can alternatively restrict to adversaries that send pairs of identical records at every step except for one. Under this formulation, the definition of differential privacy proposed by [13] can be restated as follows: A mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP differential privacy against adaptive adversaries if for every such adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously. Again, below we will enforce the restriction on the adversary via another IPM $\mathcal{V}$.
@@
  • Any non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ as a response. Upon receiving any subsequent input message, the CM representing $\mathcal{N}$ halts the communication.
  • Any interactive mechanism $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ can be modeled as a CM that receives an initial state $s\in S_{\mathcal{M}}^{\text{init}}$ as its first message and responds with an acknowledgment message $\top$. This CM then answers the next input messages the same as $\mathcal{M}(s)$.
@@
For a verification function $f$, a verifier $\mathcal{V}[f]$ is an IPM with the initial state space $S_{\mathcal{V}[f]}^{\text{init}}=\{()\}$, where ()()( ) represents an empty sequence. Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state ()()( ) as $\mathcal{V}[f]$. $\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms. The state of $\mathcal{V}[f]$ consists of the sequence of messages it has received from the left mechanism so far. Upon receiving a message $m$ from the left mechanism, $\mathcal{V}[f]$ appends $m$ to its state and applies the verification function $f$. If the sequence is $f$-valid, then $\mathcal{V}[f]$ forwards the message $m$ to the right mechanism; otherwise, it halts the communication. $\mathcal{V}[f]$ forwards any message from the right mechanism unchanged to the left mechanism.
@@
By definition, for any continual mechanism $\mathcal{M}$ and any verification function $f$, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is an IM with two possible initial states. The mechanism $\mathcal{I}\!\circ^{*}\!\mathcal{M}$ accepts messages of the form $(m_{0},m_{1})$, where both $m_{0}$ and $m_{1}$ are in the query space $Q_{\mathcal{M}}$ of $\mathcal{M}$. Therefore, when analyzing $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$, we assume that the verification function $f$ always returns $\bot$ whenever a message in its input sequence is not in $Q_{\mathcal{M}}\times Q_{\mathcal{M}}$, even if not explicitly say so.
@@
Let $\mathcal{M}$ be a CM and $f$ a verification function. We say that $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private (or $(\epsilon,\delta)$-DP for short) w.r.t. $f$ against adaptive adversaries if for any adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
[…]
@@ 4. Concurrent Composition of Continual Mechanisms
The initial state space of ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$, where $S_{\mathcal{M}_{i}}^{\text{init}}$ is the initial state space of $\mathcal{M}_{i}$. Let $\sim_{i}$ be a neighboring relation on each $S_{\mathcal{M}_{i}}^{\text{init}}$, and define the neighboring relation $\sim$ on $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$ as follows: for any two initial states $s=(s_{1},\dots,s_{k})$ and $s^{\prime}=(s_{1}^{\prime},\dots,s_{k}^{\prime})$, we have $s\sim s^{\prime}$ if and only if $s_{i}\sim_{i}s_{i}^{\prime}$ for every $i\in[k]$.
@@
To extend Lyu’s result on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and any pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent. However, as we discuss in Section 4.2, there are hidden assumptions in the proof of this lemma. Additionally, we show there that the construction of $\mathcal{P}$ in [20], for $\epsilon_{i}$-DP IMs, satisfies a property, required for a later proof.
@@
Upon receiving $(\mathcal{M},s,\epsilon,\delta,f)$, ExtConComp creates an instance of $\mathcal{M}$ with the initial state $s$, and when receiving $(q,i)$, it forwards $q$ to the $i$-th created mechanism and returns its response. If less than $i$ mechanisms have been created, ExtConComp halts the communication. The state of ExtConComp is a sequence of pairs $(\mathcal{M}_{i},s_{i}^{\prime})$, where the $i$-th pair represents the $i$-th created mechanism $\mathcal{M}_{i}$ and its current state $s_{i}^{\prime}$. The single initial state space of ExtConComp is the empty sequence ()()( ). Both query set and answer set of ExtConComp are the universal message set $\mathcal{U}$ in Remark 3.4.
@@
For any verification function $f$, the $f$-concurrent composition of CMs is the interactive mechanism $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
@@
A sequence of messages $m_{1},\dots,m_{k}$ is suitable if for every $1\leq j\leq k$, (i) $m_{j}$ is of the form $(m_{j}^{0},m_{j}^{1})$, where $m_{j}^{0}$ and $m_{j}^{1}$ are either identical creation queries or messages of the forms $(q_{0},\ell)$ and $(q_{1},\ell)$ with the same index $\ell$; (ii) for each $b\in\{0,1\}$, if $m_{j}^{b}=(q,\ell)$, then there are at least $\ell$ creation queries in the sequence $m_{1}^{b},\dots,m_{j-1}^{b}$; (iii) for each $b\in\{0,1\}$, if $m_{j}^{b}=(q,\ell)$ and $\mathcal{M}_{\ell}$ is the $\ell$-th mechanism requested to be created in $m_{1}^{b},\dots,m_{j-1}^{b}$, then $q$ is in the query space of $\mathcal{M}_{\ell}$, i.e., $q\in Q_{\mathcal{M}_{\ell}}$; and (iv) if $m_{j}^{0}=m_{j}^{1}=(\mathcal{M},s,\epsilon,\delta,f)$, then $\mathcal{M}$ is a CM with a single initial state $s$, and it is $(\epsilon,\delta)$-DP w.r.t. $f$.
@@ 4.2. $(\epsilon,\delta)$-DP IMs and $\mathrm{RR}_{\epsilon,\delta}$
Lyu [18] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are identical. Vadhan and Wang [20] show the same result in the special case where $\delta=0$.
Although not explicitly stated, the constructions of $\mathcal{T}$ proposed in both [18] and [20] require the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be discrete. This assumption is formally defined as:
@@
Lyu [18] additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers. In Appendix A, we will modify their proof to remove this additional assumption. Let $\mathcal{M}^{\prime}$ be a continual mechanism that is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$. By Corollary 3.8, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$ is an IM that is $(\epsilon,\delta)$-DP w.r.t. $\sim_{01}$. Replacing $\mathcal{M}$ in Lemma A.1 with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$, we get:
@@
For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}$ be a CM that is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$. Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions and responds to at most $c$ queries. Then, there exists an IPM $\mathcal{T}$ with a single initial state ()()( ) such that for each $b\in\{0,1\}$, the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent.
Let $\mathcal{M}$ be an IM that is $\epsilon$-DP w.r.t. a neighbor relation $\sim$. For every pair of neighboring initial states $s_{0}$ and $s_{1}$, Vadhan and Wang [20] construct an IPM $\mathcal{T}$ (with a single initial state) such that for each bit $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,0}(b)$ and $\mathcal{M}(s_{b})$ are identical. Appendix B, we will review their construction and show that for any query sequence $q_{1},\dots,q_{t}$, if the distributions of the $t$ answers of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to the queries $q_{1},\dots,q_{t}$ are identical, then $\mathcal{T}$ does not need to know the output of $\mathrm{RR}_{\epsilon,0}(b)$ in the first $t$ steps. This observation allows us to delay the interaction of $\mathcal{T}$ and $\mathrm{RR}_{\epsilon,0}(b)$ as much as possible.
Let $\mathcal{M}=\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$, where $\mathcal{M}^{\prime}$ is an $\epsilon$-DP CM w.r.t. $f$. By definition, upon receiving a query sequence $q_{1},\dots,q_{t}$ where each $q_{j}=(q_{j}^{\prime},q_{j}^{\prime})$ is a pair of identical queries, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}^{\prime}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M}^{\prime}$ have the same answer distribution, implying the following lemma:
Lemma 4.7 ([20]).
For $\epsilon\geq 0$, let $\mathcal{M}$ be a CM that is $\epsilon$-DP w.r.t. a verification function $f$. Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions. Then, there exists an IPM $\mathcal{T}$ with a single initial state ()()( ) such that for each $b\in\{0,1\}$,
  • the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,0}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent, and
  • as long as $\mathcal{T}$ receives pairs of identical queries, it does not interact with $\mathrm{RR}_{\epsilon,0}(b)$.
@@
Since $\mathrm{RR}_{\epsilon,\delta}$ is a NIM, in both Lemma 4.6 and Lemma 4.7, the IPM $\mathcal{T}$ interacts with $\mathrm{RR}_{\epsilon,\delta}$ at most once.

### v3

@@ 3. Continual Mechanisms
By Definition 2.13, an interactive mechanism $\mathcal{M}$ with initial state space $S_{\mathcal{M}}^{\text{init}}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$, if for every two initial states $s_{0},s_{1}\in S_{\mathcal{M}}^{\text{init}}$ such that $s_{0}\sim s_{1}$ and every adversary $\mathcal{A}$, the view of $\mathcal{A}$ in the interaction with $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ is $(\epsilon,\delta)$-indistinguishable. This means that for the worst-case (i.e., adversarial) choice of neighboring initial states, an adversary by adaptively asking queries should not be able to determine whether $\mathcal{M}$ is running on $s_{0}$ or $s_{1}$.
@@
To formalize this, we restrict attention to adversaries that send a pair of neighboring initial states $(s_{0},s_{1})$ as their first message and pairs of identical queries $(q_{i},q_{i})$ in its subsequent messages. We also define an IPM $\mathcal{I}$ with the initial state space $S_{\mathcal{I}}^{\text{init}}=\{0,1\}$ to stand between the adversary and $\mathcal{M}$, filtering the pairs based on its, i.e., $\mathcal{I}$’s, initial state. Specifically, for $b\in\{0,1\}$, in an interaction between an adversary $\mathcal{A}$ and $\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$, upon receiving a message $(m_{0},m_{1})$ from $\mathcal{A}$, $\mathcal{I}(b)$ forwards $m_{b}$ to $\mathcal{M}$ and returns $\mathcal{M}$’s answer unchanged to $\mathcal{A}$. With this formulation, an interactive mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for every adversary of the above form, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
@@
To capture ”event-level privacy” (where the privacy unit is a single record), Dwork et al. [7] define neighboring as a binary relation on sequences of data and empty records: two sequences are neighboring if they are identical except for the presence or absence of a record in a single step. They called a mechanism $(\epsilon,\delta)$-DP under continual observation if the distributions of answer sequences are $(\epsilon,\delta)$-indistinguishable for every two neighboring input sequences. This definition does not allow for data sequences that are adaptively chosen based on the answers released.
@@
A mechanism $\mathcal{M}$ in the continual observation model can be viewed as an IM with a single possible initial state, representing an empty dataset. The input message set of $\mathcal{M}$ consists of the empty record and all possible data records. Instead of having a challenge step, we can alternatively restrict to adversaries that send pairs of identical records at every step except for one. Under this formulation, the definition of differential privacy proposed by [14] can be restated as follows: A mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP differential privacy against adaptive adversaries if for any such adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{I}[0]\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{I}[1]\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable, where $\mathcal{I}$ is the interactive post-processing mechanism discussed previously. Again, below we will enforce the restriction on the adversary via another IPM $\mathcal{V}$.
@@
  • Every non-interactive mechanism $\mathcal{N}$, such as the Laplace mechanism, can be modeled as a CM that receives a dataset $x$ as its first message and returns $\mathcal{N}(x)$ as a response. Upon receiving any subsequent input message, the CM representing $\mathcal{N}$ halts the communication.
  • Every interactive mechanism $\mathcal{M}$ with an initial state space $S_{\mathcal{M}}^{\text{init}}$ can be modeled as a CM that receives an initial state $s\in S_{\mathcal{M}}^{\text{init}}$ as its first message and responds with an acknowledgment message $\top$. This CM then answers the next input messages the same as $\mathcal{M}(s)$.
@@
For a verification function $f$, a verifier $\mathcal{V}[f]$ is an IPM with the initial state space $S_{\mathcal{V}[f]}^{\text{init}}=\{()\}$, where $()$ represents an empty sequence. Since the initial state space of $\mathcal{V}[f]$ is a singleton, we denote an instance of $\mathcal{V}[f]$ with the initial state $()$ as $\mathcal{V}[f]$. $\mathcal{V}[f]$ stands between two IMs, referred to as the left and right mechanisms. The state of $\mathcal{V}[f]$ consists of the sequence of messages it has received from the left mechanism so far. Upon receiving a message $m$ from the left mechanism, $\mathcal{V}[f]$ appends $m$ to its state and applies the verification function $f$. If the sequence is $f$-valid, then $\mathcal{V}[f]$ forwards the message $m$ to the right mechanism; otherwise, it halts the communication. $\mathcal{V}[f]$ forwards any message from the right mechanism unchanged to the left mechanism.
@@
By definition, for every continual mechanism $\mathcal{M}$ and every verification function $f$, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is an IM with two possible initial states. The mechanism $\mathcal{I}\!\circ^{*}\!\mathcal{M}$ accepts messages of the form $(m_{0},m_{1})$, where both $m_{0}$ and $m_{1}$ are in the query space $Q_{\mathcal{M}}$ of $\mathcal{M}$. Therefore, when analyzing $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$, we assume that the verification function $f$ always returns $\bot$ whenever a message in its input sequence is not in $Q_{\mathcal{M}}\times Q_{\mathcal{M}}$, even if not explicitly say so.
@@
Let $\mathcal{M}$ be a CM and $f$ a verification function. We say that $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private (or $(\epsilon,\delta)$-DP for short) w.r.t. $f$ against adaptive adversaries if for every adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
Let $S_{\mathcal{M}}^{\text{init}}=\{s^{\text{init}}\}$ be the initial state space of the continual mechanism $\mathcal{M}$ in Definition 3.7. By definition, the initial state spaces of $\mathcal{I}$ and $\mathcal{V}[f]$ are $\{0,1\}$ and $\{()\}$, respectively. Thus, the initial state space of $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ is $\left\{((),0,s^{\text{init}}),((),1,s^{\text{init}})\right\}$. Define $\sim_{01}$ as a neighboring on this set, where only $((),0,s^{\text{init}})\sim_{01}((),1,s^{\text{init}})$. Comparing Definition 3.7 and Definition 2.13 leads to the following corollary:
@@ 4. Concurrent Composition of Continual Mechanisms
The initial state space of ${\textrm{ConComp}}[\mathcal{M}_{1},\dots,\mathcal{M}_{k}]$ is $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$, where $S_{\mathcal{M}_{i}}^{\text{init}}$ is the initial state space of $\mathcal{M}_{i}$. Let $\sim_{i}$ be a neighboring relation on each $S_{\mathcal{M}_{i}}^{\text{init}}$, and define the neighboring relation $\sim$ on $S_{\mathcal{M}_{1}}^{\text{init}}\times\dots\times S_{\mathcal{M}_{k}}^{\text{init}}$ as follows: for every two initial states $s=(s_{1},\dots,s_{k})$ and $s^{\prime}=(s_{1}^{\prime},\dots,s_{k}^{\prime})$, we have $s\sim s^{\prime}$ if and only if $s_{i}\sim_{i}s_{i}^{\prime}$ for all $i\in[k]$.
@@
To extend Lyu’s result on the privacy guarantee of the concurrent composition of IMs to the concurrent composition of CMs, we want to use the following intermediate lemma from their paper: For every $(\epsilon,\delta)$-DP interactive mechanism $\mathcal{M}$ and every pair of neighboring initial states $s$ and $s^{\prime}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the IMs $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(b)$ are equivalent. However, as we discuss in Section 4.2, there are hidden assumptions in the proof of this lemma. Additionally, we show there that the construction of $\mathcal{P}$ in [23], for $\epsilon_{i}$-DP IMs, satisfies a property, required for a later proof.
@@
Upon receiving $(\mathcal{M},s,\epsilon,\delta,f)$, ExtConComp creates an instance of $\mathcal{M}$ with the initial state $s$, and when receiving $(q,i)$, it forwards $q$ to the $i$-th created mechanism and returns its response. If less than $i$ mechanisms have been created, ExtConComp halts the communication. The state of ExtConComp is a sequence of pairs $(\mathcal{M}_{i},s_{i}^{\prime})$, where the $i$-th pair represents the $i$-th created mechanism $\mathcal{M}_{i}$ and its current state $s_{i}^{\prime}$. The single initial state space of ExtConComp is the empty sequence $()$. Both query set and answer set of ExtConComp are the universal message set $\mathcal{U}$ in Remark 3.4.
@@
Let $f$ be a verification function. The $f$-concurrent composition of CMs is the interactive mechanism $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$.
@@
A sequence of messages $m_{1},\dots,m_{k}$ is said to be suitable if the following conditions hold for every $1\leq j\leq k$:
  1. (i)
    Each message $m_{j}$ is either
    1. (a)
      a pair of identical creation queries of the form $(\mathcal{M},s,\epsilon,\delta,f)^{2}$, or
    2. (b)
      a query pair message of the form $\left((q_{0},\ell),(q_{1},\ell)\right)$, where both queries $q_{0}$ and $q_{1}$ share the same index $\ell$.
  2. (ii)
    If $m_{j}=\left((q_{0},\ell),(q_{1},\ell)\right)$, then there are at least $\ell$ creation-query pairs in the sequence $m_{1},\dots,m_{j-1}$.
  3. (iii)
    If $m_{j}=\left((q_{0},\ell),(q_{1},\ell)\right)$ and $\mathcal{M}_{\ell}$ is the $\ell$-th mechanism requested to be created in the preceding messages $m_{1},\dots,m_{j-1}$, then both $q_{0}$ and $q_{1}$ belong to the query space of $\mathcal{M}_{\ell}$, i.e., $q_{0},q_{1}\in Q_{\mathcal{M}_{\ell}}$.
  4. (iv)
    If $m_{j}=(\mathcal{M},s,\epsilon,\delta,f)^{2}$, then $\mathcal{M}$ is a CM with a single initial state $s$, and it is $(\epsilon,\delta)$-DP w.r.t. $f$.
  5. (v)
    If $m_{j}=(\mathcal{M},s,\epsilon,\delta,f)^{2}$ is the $\ell$-th creation-query pair in the sequence $m_{1},\dots,m_{j-1}$, then the corresponding sequence of query pairs $(q_{0},q_{1})$ from the messages of the form $\left((q_{0},\ell),(q_{1},\ell)\right)$ in $m_{1},\dots,m_{k}$ is $f$-valid.
@@ 4.2. $(\epsilon,\delta)$-DP IMs and $\mathrm{RR}_{\epsilon,\delta}$
Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{T}$ (with a single initial state) such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{M}(s_{b})$ are identical. Although not explicitly stated, the construction of $\mathcal{T}$ proposed in [19] requires the answer distributions of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to be discrete. This assumption is formally defined as:
@@
Lyu [19] additionally assumes that there exists a pre-fixed upper bound $c$ on the number of queries $\mathcal{M}$ answers. This assumption stems from their use of a backward-defined control function. By employing the limit control functions which we introduce in Section 8, one can modify their proof to remove this additional assumption. Moreover, using the formalization of Section 8 to define recursive functions, we can relax the requirement that the answer set is finite to allow countably infinite answer sets.
Let $\mathcal{M}^{\prime}$ be a continual mechanism that is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$. By Corollary 3.8, $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}^{\prime}$ is an IM that is $(\epsilon,\delta)$-DP w.r.t. $\sim_{01}$. We have:
@@
For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}$ be a CM that is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$. Suppose $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!\mathcal{M}$ has discrete answer distributions. Then, there exists an IPM $\mathcal{T}$ with a single initial state such that for each $b\in\{0,1\}$, the IMs $\mathcal{T}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}$ are equivalent.
@@
Since $\mathrm{RR}_{\epsilon,\delta}$ is a NIM, in Lemma 4.6, the IPM $\mathcal{T}$ interacts with $\mathrm{RR}_{\epsilon,\delta}$ at most once.